# Reservoir-assisted quantum battery charging at finite temperatures

### v2

@@ IV MULTIPARTICLE QUANTUM BATTERY MODEL
On the basis of the quantum feedback mechanism, we extend the previously proposed single-particle battery model to a multi-particle configuration. In this extended model, the charger consists of a single two-level atom embedded in a zero-temperature reservoir. During the charging process, quantum feedback control is applied to the charger. The quantum battery is composed of multiple two-level atoms and is embedded in a common thermal reservoir. The coupling strength between the charger and each atom in the battery is $g$. The free Hamiltonian form of the multi-particle quantum battery is given by form
@@
where $S_{B}^{x,y,z}=\sum_{i}\sigma_{iB}^{x,y,z}/2$ denotes the collective atom operators of the quantum battery and $N$ denotes the number of atoms comprising the battery.
In the Markovian limit (i.e.,$\tau\rightarrow 0$), the evolution of the system is described by the following master equation
@@
At this point, the interaction term $H_{\rm SI}^{{}^{\prime}}$ between the charger and the quantum battery, as well as the Lindblad term $\mathcal{L}_{B}^{\prime}\rho$ for the quantum battery, is given by the following forms
$H_{\rm SI}^{\prime}=\sum_{i=1}^{N}g_{i}(\sigma_{C}^{+}\sigma_{iB}^{-}+\sigma_{C}^{-}\sigma_{iB}^{+})+\sum_{i<j}^{N}J_{i}(\sigma_{iB}^{+}\sigma_{jB}^{-}+\sigma_{iB}^{-}\sigma_{jB}^{+}),$ (23)
@@
$+\gamma^{\uparrow}_{B}[L_{B}^{+}\rho L_{B}^{-}-1/2(L_{B}^{-}L_{B}^{+}\rho+\rho L_{B}^{-}L_{B}^{+})],$
the first term of Eq. (23) describes the direct coupling between the charger and the constituent atoms of the quantum battery, while the second term accounts for the interactions between the atoms within the battery itself. For simplicity, we assume $g_{i}=g$ and $J_{i}=J$. In Eq. (24), $L_{B}^{-}=\sum_{i=1}^{N}\sigma_{iB}^{-}$ denotes the collective lowering operator of the quantum battery. To facilitate a fair comparison with the single-particle quantum battery, the steady state energy density of the multi-particle quantum battery is defined as follows
$\mathcal{W}=\frac{E_{B}(\infty)}{N}=\omega_{0}\left(\frac{\langle S_{B}^{z}(\infty)\rangle}{N}+\frac{1}{2}\right),$ (25)
@@
Figure 15: In the bosonic thermal reservoir, panels (a)-(c) depict the variation of the average ergotropy of the quantum battery in the steady state as a function of the particle number $N$, with the corresponding atom-atom coupling strengths in the multi-particle quantum battery set to $J=0$, $J=0.1g$, and $J=g$, respectively. Panels (d)-(f) represent the optimal dissipation rate $\Gamma_{C}$ of the charger corresponding to panels (a)-(c). Different curves represent different temperatures, while the other parameters are identical to those in Fig. 14.
@@
Figure 16: The multi-particle quantum battery is placed in a bosonic thermal reservoir. The charging efficiency $R$ of the battery is analyzed as a function of the number of particles $N$. The interparticle interaction strengths of the multi-particle quantum battery in panels (a)-(c) are set to $J=0$, $J=0.1g$, and $J=g$, respectively. Different curves represent different temperatures, while all other parameters are identical to those in Fig. 14.
@@
In the following, we investigate the performance of the multi-particle quantum battery under different thermal reservoirs. The feedback parameter is set to $\delta=f/\Gamma_{C}=1$ to facilitate a more meaningful comparison with the single-particle quantum battery in the subsequent analysis.
@@
IV.1 The charging process of the multi-particle quantum battery in a bosonic thermal reservoir
@@
First, we investigate the charging process of the multi-particle quantum battery in a bosonic thermal reservoir, in this case, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1+n_{b})$ and $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{b}$ in Eq. (24). To explore the effects of the number of quantum battery particles $N$ and the interparticle interactions $J$ on battery performance, we examine the optimal energy density of the battery as a function of $N$ under three coupling regimes: no coupling ($J=0$), weak coupling ($J=0.1g$) and strong coupling ($J=g$), as shown in Figs. 14 (a)-(c). The quantum battery operates in the desired weak dissipation regime with $\Gamma_{B}=0.05g$, and the different curves correspond to different thermal reservoir temperatures. Additionally, the corresponding optimal charger dissipation rates for Figs. 14 (a)-(c) are presented in Figs. 14 (d)-(f). The results indicate that, similar to the single-particle quantum battery case, temperature remains a detrimental factor for charging. Furthermore, it is observed that, under low-temperature conditions, the energy density of the battery increases with the number of particles when the interparticle couplings are absent or weak, as shown in Fig. 14 (a) and Fig. 14 (b). However, at higher temperatures, the energy density decreases as particle number increases, thus impairing efficient energy transfer to the battery. Furthermore, in the strong interparticle coupling regime, both elevated temperature and an increased number of particles have detrimental effects on the energy storage capacity of the battery, as illustrated in Fig. 14(c).
@@
Figure 17: The multi-particle quantum battery is placed in a bosonic thermal reservoir. The solid line represents the energy density of the multi-particle quantum battery, while the dashed line represents the maximum extractable energy per particle (average ergotropy). Curves of different colors correspond to different temperatures, with the red curve representing $T=0$ ($n_{b}=0$) and the blue curve representing $T=10$ [$n_{b}=1/(e^{1/10}-1)\approx 9.51$]. Panels (a) and (b) correspond to 3-particle quantum batteries with dissipation rates $\Gamma_{B}=0.05g$ and $\Gamma_{B}=0.5g$, respectively, while Panels (c) and (d) correspond to 5-particle quantum batteries with the same dissipation rates. All other parameters are identical to those in Fig. 14.
@@
Figure 18: The multi-particle quantum battery is placed in a fermionic thermal reservoir. Panels (a)-(c) depict the variation in the steady-state energy density of the multi-particle quantum battery as a function of particle number $N$ for $J=0$, $J=0.1g$, and $J=g$, respectively. Panels (d)-(f) correspond to the optimal charger dissipation parameters required for Panels (a)-(c). Different curves correspond to different temperatures. The other parameters are the same as those in Fig. 14.
@@
Subsequently, Fig. 15 presents the dependence of the maximum extractable energy per particle (i.e., the average ergotropy) of the multi-particle quantum battery on the number of particles $N$. As shown in Figs. 15 (a)-(b), when there is no interparticle interaction ($J=0$) or the interaction is weak ($J=0.1g$), the average ergotropy of the battery increases as the number of particles $N$ increases. Meanwhile, when the number of particles in a multi-particle quantum battery is fixed, the average ergotropy decreases as the temperature increases, which means that temperature inhibits the extraction of energy from the battery. In addition, when the interaction between particles in a multi-particle quantum battery is strong ($J=g$), the average ergotropy of the battery exhibits a slight decrease with increasing particle number $N$ at low temperatures. In contrast, at high temperatures, this behavior is suppressed, and the average ergotropy increases as the number of particles $N$ increases, but remains lower than that observed at low temperatures. As seen in Figs. 14 (d)-(f) and Figs. 15 (d)-(f), both in terms of energy storage and energy extraction, the optimal dissipation rate of the charger required increases with the number of particles. Figure 16 investigates the effects of temperature and particle number $N$ on the charging efficiency $R$ of the battery. Panels (a)–(c) correspond to $J=0$, $J=0.1g$, and $J=g$, respectively, and the different curves represent different temperatures of the thermal reservoir. It is observed that at high temperatures, the charging efficiency $R$ is more significantly affected by the number of battery particles, whereas this influence is greatly reduced at low temperatures. Overall, temperature remains a detrimental factor for both energy storage and extraction in the multi-particle quantum battery. However, there exist two specific scenarios in which an increase in the number of quantum battery particles can be beneficial for energy extraction: one occurs with weak interparticle interactions, and the other at high temperatures with strong interparticle interactions. The multi-particle quantum battery operates in a strong dissipation regime, as discussed in the Appendix C.
@@
In the following, we further investigate the effect of interparticle interaction strength $J$ on the energy density and the average ergotropy of multi-particle quantum batteries by analyzing the cases of three-particle and five-particle quantum batteries. Figure 17 plots the variation of the energy density (solid lines) and the average ergotropy (dashed lines) with interparticle interaction $J$ at different temperatures and the dissipation rates of the battery $\Gamma_{B}$. Specifically, Figs. 17 (a)-(b) correspond to the case of the 3-particle quantum battery, with dissipation rates of $\Gamma_{B}=0.05g$ and $\Gamma_{B}=0.5g$, respectively. Figures 17 (c)-(d) correspond to the case of the 5-particle quantum battery. The curves in different colors represent different temperatures of the bosonic thermal reservoir, with the red curve corresponding to $T=0$ and the blue curve corresponding to $T=10$. The results indicate that, regardless of whether the quantum battery operates under a weak or strong dissipation regime, interparticle interactions tend to degrade the performance of the multi-particle quantum battery. Compared to the weak dissipation regime, strong dissipation significantly impairs the energy storage and extraction capabilities of the quantum battery. However, under a strong dissipation regime, a high-temperature environment can enhance both energy storage and extraction.
@@
Figure 19: The multi-particle quantum battery is embedded in a fermionic thermal reservoir. The average ergotropy of the multi-particle quantum battery is plotted as a function of the number of particles $N$, with all other parameters consistent with those in Fig. 14.
@@
Figure 20: Panels (a)-(c) respectively depict the variation in the charging efficiency $R$ of the multi-particle quantum battery in the steady state with the number of particles $N$ in a fermionic thermal reservoir. The other parameters are the same as those in Fig. 14.
@@
Figure 21: Similar to Fig. 17, panels (a)-(b) and (c)-(d) respectively show the variations in the energy density and the average ergotropy for the 3-particle and 5-particle quantum batteries under weak dissipation rate $\Gamma_{B}=0.05g$ and strong dissipation rate $\Gamma_{B}=0.5g$, as functions of the interparticle coupling strength $J$ of the battery. The red curve corresponds to $T=0$ ($n_{f}=0$) and the blue curve corresponds to $T=10$ [$n_{f}=1/(e^{1/10}+1)\approx 0.475$].
@@
IV.2 The charging process of the multi-particle quantum battery in a fermionic thermal reservoir
@@
Next, we investigate the charging process of the multi-particle quantum battery coupled to a fermionic thermal reservoir. In this scenario, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1-n_{f})$ and $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{f}$ in Eq. (24).
@@
Firstly, we explore the effects of a finite-temperature fermionic thermal reservoir on the energy storage and maximum extractable energy (ergotropy) of the multi-particle quantum battery, under the weak dissipation condition ($\Gamma_{B}=0.05g$). Furthermore, we examine the effects of the interparticle interaction strength $J$ and the number of particles $N$ on the battery performance, aiming to identify a more efficient charging mechanism.
@@
Figures 18 (a)-(c) respectively illustrate the variation of energy density (i.e., the stored energy per particle) of the multi-particle quantum battery as a function of the number of particles $N$, under three different conditions: no coupling ($J=0$), weak coupling ($J=0.1g$), and strong coupling ($J=g$) between the internal particles of the battery. The optimal charger dissipation rates corresponding to Figs. 18 (a)-(c) are presented in Figs. 18 (d)-(f). The numerical results show that when the interaction strength $J$ between particles is relatively weak ($J=0$ or $0.1g$), increasing the number of battery particles enhances the energy density of the multi-particle quantum battery. Moreover, a higher temperature also leads to an increase in energy density. Similar to the single-particle battery case, temperature remains a favorable factor in the charging process. In the case of a relatively strong interaction strength $J$, it can be observed from Fig. 18 (c) that the energy density of the multi-particle quantum battery decreases as the number of particles $N$ increases under low temperature conditions. As the temperature increases, the impact of the number of particles on energy storage becomes less pronounced. Furthermore, Figs. 18 (d)-(f) indicate that the optimal dissipation rate $\Gamma_{C}$ of the charger required to maximize energy storage increases with the number of battery particles $N$ increases.
@@
Figure 19 illustrates the impact of the reservoir temperature, particle number $N$, and the interparticle interaction $J$ on the average ergotropy of the multi-particle quantum battery in a finite-temperature fermionic thermal reservoir. The results indicate that the behavior of the ergotropy closely resembles that of the energy storage in the battery. Figure 20 reveals the influences of the temperature and the particle number $N$ on the charging efficiency $R$ of the multi-particle quantum battery. First, it can be observed that an increase in temperature enhances the charging efficiency $R$ of the multi-particle quantum battery. Second, when the interparticle coupling strength is weak ($J=0$ or $0.1g$), $R$ increases as the particle number $N$ increases. However, when the interparticle coupling strength $J$ is strong ($J=g$), an increase in $N$ leads to a reduction in the charging efficiency $R$ at low temperatures. In contrast, at high temperatures, the effect of increasing particle number on the charging efficiency becomes less pronounced. Based on the above research findings, we observe that when the dissipation rate of the multi-particle quantum battery in a fermionic thermal reservoir is weak, higher temperatures and weaker interparticle coupling enhance the performance of the quantum battery. The case of a strongly dissipative quantum battery is discussed in the Appendix D.
@@
Subsequently, we continue to examine the impact of interparticle interaction $J$ on both the energy density and average ergotropy of the multi-particle quantum battery, using 3-particle and 5-particle quantum batteries as representative cases, as shown in Fig. 21. Panels (a)-(b) display the numerical results for the three-particle quantum battery with dissipation rates $\Gamma_{B}=0.05g$ and $\Gamma_{B}=0.5g$, respectively, while panels (c)-(d) show the results for corresponding five-particle quantum battery under identical dissipation conditions. The red and blue curves correspond to the fermionic thermal reservoir temperatures of the $T=0$ and $T=10$, respectively. The solid lines represent the energy density of the battery, while the dashed lines represent the average ergotropy. Numerical analysis reveals that when the multi-particle quantum battery is coupled to a fermionic thermal reservoir, interparticle interactions $J$ do not enhance its performance, regardless of whether the battery operates in a weak or strong dissipation regime. However, compared to low-temperature conditions, a high-temperature environment significantly improves both the energy storage capacity and the extractable energy of the battery.
@@
Based on the above discussion, regardless of whether the multi-particle quantum battery is placed in a bosonic or fermionic thermal reservoir, both the energy density and the average ergotropy of the multi-particle quantum battery decrease as the interparticle interaction strength $J$ increases. The underlying physical mechanism behind this behavior lies in the interaction-induced splitting of the battery’s energy eigenstates. This spectral reconstruction shifts the system away from the resonance, thereby suppressing energy transfer from the charger to the quantum battery. Consequently, interparticle interactions within the battery emerge as a significant obstacle in designing and implementing efficient multi-particle quantum batteries. Additionally, a high-temperature fermionic thermal reservoir provides a stable foundation for building multi-particle quantum batteries.

### v3

@@ IV MULTIPARTICLE QUANTUM BATTERY MODEL
On the basis of the quantum feedback mechanism, we extend the previously proposed single-particle battery model to a multiparticle configuration. In this extended model, the charger consists of a single two-level atom embedded in a zero-temperature reservoir. During the charging process, quantum feedback control is applied to the charger. The quantum battery is composed of multiple two-level atoms and is embedded in a common thermal reservoir. The coupling strength between the charger and each atom in the battery is $g$. The free Hamiltonian form of the multiparticle quantum battery is given as
@@
where $S_{B}^{x,y,z}=\sum_{i}\sigma_{iB}^{x,y,z}/2$ defines the collective atom operators of the quantum battery and $N$ denotes the number of atoms comprising the battery.
In the Markovian limit (i.e.,$\tau\rightarrow 0$), the evolution of the system is described by the master equation
@@
At this point, the interaction term $H_{SI}^{\prime}$ between the charger and the quantum battery, and the Lindblad term $\mathcal{L}_{B}^{\prime}\rho$ for the quantum battery, are given in the forms
$H_{SI}^{\prime}=\sum_{i=1}^{N}g_{i}(\sigma_{C}^{+}\sigma_{iB}^{-}+\sigma_{C}^{-}\sigma_{iB}^{+})+\sum_{i<j}^{N}J_{i}(\sigma_{iB}^{+}\sigma_{jB}^{-}+\sigma_{iB}^{-}\sigma_{jB}^{+}).$ (23)
@@
$+\gamma^{\uparrow}_{B}[L_{B}^{+}\rho L_{B}^{-}-1/2(L_{B}^{-}L_{B}^{+}\rho+\rho L_{B}^{-}L_{B}^{+})].$
The first term of Eq. (23) describes the direct coupling between the charger and the constituent atoms of the quantum battery, while the second term accounts for the interactions between the atoms within the battery itself. For simplicity, we assume $g_{i}=g$ and $J_{i}=J$. In Eq. (24), $L_{B}^{-}=\sum_{i=1}^{N}\sigma_{iB}^{-}$ defines the collective lowering operator of the quantum battery. To facilitate a fair comparison with the single-particle quantum battery, the steady-state energy density of the multiparticle quantum battery is defined as
$\mathcal{W}=\frac{E_{B}(\infty)}{N}=\omega_{0}\left(\frac{\langle S_{B}^{z}(\infty)\rangle}{N}+\frac{1}{2}\right).$ (25)
@@
Figure 15: In the bosonic thermal reservoir, the variation of the average ergotropy of the quantum battery in the steady state is plotted as a function of the particle number $N$, with the corresponding atom-atom coupling strengths in the multiparticle quantum battery set to (a) $J=0$, (b) $J=0.1g$, and (c) $J=g$. (d)-(f) Optimal dissipation rate $\Gamma_{C}$ of the charger corresponding to (a)-(c). Different curves represent different temperatures, the other parameters are identical to those in Fig. 14.
@@
Figure 16: The multiparticle quantum battery is placed in a bosonic thermal reservoir. The charging efficiency $R$ of the battery is analyzed as a function of the number of particles $N$. The interparticle interaction strengths of the multiparticle quantum battery are set to (a) $J=0$, (b) $J=0.1g$, and (c) $J=g$. Different curves represent different temperatures; all other parameters are identical to those in Fig. 14.
@@
In the following, we investigate the performance of the multiparticle quantum battery under different thermal reservoirs. The feedback parameter is set to $\delta=f/\Gamma_{C}=1$ to facilitate a more meaningful comparison with the single-particle quantum battery in the subsequent analysis.
@@
IV.1 Charging process of the multiparticle quantum battery in a bosonic thermal reservoir
@@
First, we investigate the charging process of the multiparticle quantum battery in a bosonic thermal reservoir, in this case, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1+n_{b})$ and $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{b}$ in Eq. (24). To explore the effects of the number of quantum battery particles $N$ and the interparticle interactions $J$ on battery performance, we examine the optimal energy density of the battery as a function of $N$ under three coupling regimes: no coupling ($J=0$), weak coupling ($J=0.1g$), and strong coupling ($J=g$), as shown in Figs. 14(a)-14(c). The quantum battery operates in the desired weak dissipation regime with $\Gamma_{B}=0.05g$, and the different curves correspond to different thermal reservoir temperatures. Additionally, the corresponding optimal charger dissipation rates for Figs. 14(a)-14(c) are presented in Figs. 14(d)-14(f). The results indicate that, similar to the single-particle quantum battery case, temperature remains a detrimental factor for charging. Furthermore, it is observed that, under low-temperature conditions, the energy density of the battery increases with the number of particles when the interparticle couplings are absent or weak, as shown in Figs. 14(a) and 14(b). However, at higher temperatures, the energy density decreases as particle number increases, thus impairing efficient energy transfer to the battery. Furthermore, in the strong interparticle coupling regime, both elevated temperature and an increased number of particles have detrimental effects on the energy storage capacity of the battery, as illustrated in Fig. 14(c).
@@
Figure 17: The multiparticle quantum battery is placed in a bosonic thermal reservoir. The solid line represents the energy density of the multiparticle quantum battery, while the dashed line represents the maximum extractable energy per particle (average ergotropy). Curves of different colors correspond to different temperatures, with the red curve representing $T=0$ ($n_{b}=0$) and the blue curve representing $T=10$ [$n_{b}=1/(e^{1/10}-1)\approx 9.51$]. Plotted are three-particle quantum batteries with dissipation rates (a) $\Gamma_{B}=0.05g$ and (b) $\Gamma_{B}=0.5g$ and five-particle quantum batteries with dissipation rates (c) $\Gamma_{B}=0.05g$ and (d) $\Gamma_{B}=0.5g$. All other parameters are identical to those in Fig. 14.
@@
Figure 18: The multiparticle quantum battery is placed in a fermionic thermal reservoir. The depict the variation in the steady-state energy density of the multiparticle quantum battery is plotted as a function of particle number $N$ for (a) $J=0$, (b) $J=0.1g$, and (c) $J=g$. (d)-(f) Optimal charger dissipation parameters required for (a)-(c). Different curves correspond to different temperatures. The other parameters are the same as those in Fig. 14.
@@
Subsequently, Fig. 15 presents the dependence of the maximum extractable energy per particle (i.e., the average ergotropy) of the multiparticle quantum battery on the number of particles $N$. As shown in Figs. 15(a)-15(b), when there is no interparticle interaction ($J=0$) or the interaction is weak ($J=0.1g$), the average ergotropy of the battery increases as the number of particles $N$ increases. Meanwhile, when the number of particles in a multiparticle quantum battery is fixed, the average ergotropy decreases as the temperature increases, which means that temperature inhibits the extraction of energy from the battery. In addition, when the interaction between particles in a multiparticle quantum battery is strong ($J=g$), the average ergotropy of the battery exhibits a slight decrease with increasing particle number $N$ at low temperatures. In contrast, at high temperatures, this behavior is suppressed, and the average ergotropy increases as the number of particles $N$ increases, but remains lower than that observed at low temperatures. As seen in Figs. 14(d)-14(f) and 15(d)-15(f), in terms of both energy storage and energy extraction, the optimal dissipation rate of the charger required increases with the number of particles. Figure 16 investigates the effects of temperature and particle number $N$ on the charging efficiency $R$ of the battery. Figures 16(a)–16(c) correspond to $J=0$, $0.1g$, and $g$, respectively, and the different curves represent different temperatures of the thermal reservoir. It is observed that at high temperatures, the charging efficiency $R$ is more significantly affected by the number of battery particles, whereas this influence is greatly reduced at low temperatures. Overall, temperature remains a detrimental factor for both energy storage and extraction in the multiparticle quantum battery. However, there exist two specific scenarios in which an increase in the number of quantum battery particles can be beneficial for energy extraction: One occurs with weak interparticle interactions, and the other at high temperatures with strong interparticle interactions. The multiparticle quantum battery operates in a strong dissipation regime, as discussed in the Appendix C.
@@
In the following, we further investigate the effect of interparticle interaction strength $J$ on the energy density and the average ergotropy of multiparticle quantum batteries by analyzing the cases of three-particle and five-particle quantum batteries. Figure 17 plots the variation of the energy density (solid lines) and the average ergotropy (dashed lines) with interparticle interaction $J$ at different temperatures and the dissipation rates of the battery $\Gamma_{B}$. Specifically, Figs. 17(a)-17(b) correspond to the case of the three-particle quantum battery, with dissipation rates of $\Gamma_{B}=0.05g$ and $\Gamma_{B}=0.5g$, respectively. Figures 17(c)-17(d) correspond to the case of the five-particle quantum battery. The curves in different colors represent different temperatures of the bosonic thermal reservoir, with the red curve corresponding to $T=0$ and the blue curve corresponding to $T=10$. The results indicate that, regardless of whether the quantum battery operates under a weak or strong dissipation regime, interparticle interactions tend to degrade the performance of the multiparticle quantum battery. Compared to the weak dissipation regime, strong dissipation significantly impairs the energy storage and extraction capabilities of the quantum battery. However, under a strong dissipation regime, a high-temperature environment can enhance both energy storage and extraction.
@@
Figure 19: The multiparticle quantum battery is embedded in a fermionic thermal reservoir. The average ergotropy of the multiparticle quantum battery is plotted as a function of the number of particles $N$, with all other parameters consistent with those in Fig. 14.
@@
Figure 20: Variation in the charging efficiency $R$ of the multiparticle quantum battery in the steady state with the number of particles $N$ in a fermionic thermal reservoir for (a) $J=0$, (b) $J=0.1g$,and (c) $J=g$. The other parameters are the same as those in Fig. 14.
@@
Figure 21: Similar to Fig. 17, variations in the energy density and the average ergotropy are shown for the (a) and (b) three-particle and (c) and (d) five-particle quantum batteries under (a) and (c) weak dissipation rate $\Gamma_{B}=0.05g$ and (b) and (d) strong dissipation rate $\Gamma_{B}=0.5g$, as a function of the interparticle coupling strength $J$ of the battery. The red curve corresponds to $T=0$ ($n_{f}=0$) and the blue curve corresponds to $T=10$ [$n_{f}=1/(e^{1/10}+1)\approx 0.475$].
@@
IV.2 Charging process of the multiparticle quantum battery in a fermionic thermal reservoir
@@
Next we investigate the charging process of the multiparticle quantum battery coupled to a fermionic thermal reservoir. In this scenario, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1-n_{f})$ and $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{f}$ in Eq. (24).
@@
First, we explore the effects of a finite-temperature fermionic thermal reservoir on the energy storage and maximum extractable energy (ergotropy) of the multiparticle quantum battery, under the weak dissipation condition ($\Gamma_{B}=0.05g$). Furthermore, we examine the effects of the interparticle interaction strength $J$ and the number of particles $N$ on the battery performance, aiming to identify a more efficient charging mechanism.
@@
Figures 18(a)-18(c) illustrate the variation of energy density (i.e., the stored energy per particle) of the multiparticle quantum battery as a function of the number of particles $N$, under three different conditions, respectively: no coupling ($J=0$), weak coupling ($J=0.1g$), and strong coupling ($J=g$) between the internal particles of the battery. The optimal charger dissipation rates corresponding to Figs. 18(a)-18(c) are presented in Figs. 18(d)-18(f). The numerical results show that when the interaction strength $J$ between particles is relatively weak ($J=0$ or $0.1g$), increasing the number of battery particles enhances the energy density of the multiparticle quantum battery. Moreover, a higher temperature also leads to an increase in energy density. Similar to the single-particle battery case, temperature remains a favorable factor in the charging process. In the case of a relatively strong interaction strength $J$, it can be observed from Fig. 18 (c) that the energy density of the multiparticle quantum battery decreases as the number of particles $N$ increases under low-temperature conditions. As the temperature increases, the impact of the number of particles on energy storage becomes less pronounced. Furthermore, Figs. 18(d)-18(f) indicate that the optimal dissipation rate $\Gamma_{C}$ of the charger required to maximize energy storage increases when the number of battery particles $N$ increases.
@@
Figure 19 illustrates the impact of the reservoir temperature, particle number $N$, and the interparticle interaction $J$ on the average ergotropy of the multiparticle quantum battery in a finite-temperature fermionic thermal reservoir. The results indicate that the behavior of the ergotropy closely resembles that of the energy storage in the battery. Figure 20 reveals the influences of the temperature and the particle number $N$ on the charging efficiency $R$ of the multiparticle quantum battery. First, it can be observed that an increase in temperature enhances the charging efficiency $R$ of the multiparticle quantum battery. Second, when the interparticle coupling strength is weak ($J=0$ or $0.1g$), $R$ increases as the particle number $N$ increases. However, when the interparticle coupling strength $J$ is strong ($J=g$), an increase in $N$ leads to a reduction in the charging efficiency $R$ at low temperatures. In contrast, at high temperatures, the effect of increasing particle number on the charging efficiency becomes less pronounced. Based on the above research findings, we observe that when the dissipation rate of the multiparticle quantum battery in a fermionic thermal reservoir is weak, higher temperatures and weaker interparticle coupling enhance the performance of the quantum battery. The case of a strongly dissipative quantum battery is discussed in the Appendix D.
@@
Subsequently, we continue to examine the impact of interparticle interaction $J$ on both the energy density and average ergotropy of the multiparticle quantum battery, using three-particle and five-particle quantum batteries as representative cases, as shown in Fig. 21. Figures 21(a)-21(b) display the numerical results for the three-particle quantum battery with dissipation rates $\Gamma_{B}=0.05g$ and $\Gamma_{B}=0.5g$, respectively, while Figs. 21(c)-21(d) show the results for the corresponding five-particle quantum battery under identical dissipation conditions. The red and blue curves correspond to the fermionic thermal reservoir temperatures of the $T=0$ and $10$, respectively. The solid lines represent the energy density of the battery, while the dashed lines represent the average ergotropy. Numerical analysis reveals that when the multiparticle quantum battery is coupled to a fermionic thermal reservoir, interparticle interactions $J$ do not enhance its performance, regardless of whether the battery operates in a weak or strong dissipation regime. However, compared to low-temperature conditions, a high-temperature environment significantly improves both the energy storage capacity and the extractable energy of the battery.
@@
Based on the above discussion, regardless of whether the multiparticle quantum battery is placed in a bosonic or fermionic thermal reservoir, both the energy density and the average ergotropy of the multiparticle quantum battery decrease as the interparticle interaction strength $J$ increases. The underlying physical mechanism behind this behavior lies in the interaction-induced splitting of the battery’s energy eigenstates. This spectral reconstruction shifts the system away from the resonance, thereby suppressing energy transfer from the charger to the quantum battery. Consequently, interparticle interactions within the battery emerge as a significant obstacle in designing and implementing efficient multiparticle quantum batteries. Additionally, a high-temperature fermionic thermal reservoir provides a stable foundation for building multiparticle quantum batteries.